A bank offers your firm a revolving credit arrangement for up to $66 million at an interest rate of 1.65 percent per quarter. The bank also requires you to maintain a compensating balance of 2 percent against the unused portion of the credit line, to be deposited in a non-interest-bearing account. Assume you have a short-term investment account at the bank that pays 1.00 percent per quarter, and assume that the bank uses compound interest on its revolving credit loans.

a. What is your effective annual interest rate (an opportunity cost) on the revolving credit arrangement if your firm does not use it during the year? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Effective annual interest rate %

b. What is your effective annual interest rate on the lending arrangement if you borrow $34 million immediately and repay it in one year? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Effective annual interest rate %

c. What is your effective annual interest rate if you borrow $66 million immediately and repay it in one year? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Effective annual interest rate %
Answer #1

Anscien Page No 0 Ebbiective anny al nieren The ebbective annual interest rate ( an opportunity corf) on the orevolving coeditheofice anual interest rate Page No ③ Effective annual interest vrate on the lending arrange -ment of you boorow $ 34 millioPage No ③ 640,000 34,000,000 = 0.01882 - 882 Opportuniling cost on non-interest becoing account - 1.882%.x4.06% = 0.000764092- Page No 4 Effective annual interest orate = (1+i)?-1 = (1+165%)4-1 = (1.0165) -- 1 = 1:0676 5 -1 = 0.06765 = 6.765 % 2 6.8